OpenStudy (anonymous):

For a triangle with sides a = 620 and b = 1050, side c must be less than what length?

OpenStudy (anonymous):

less than the sum of the other 2 sides

OpenStudy (saranya):

Assuming you are referring to a right angle triangle who's hypotenuse is c. Use the following equation: a^2 + b^2 = c^2 C = 1219.4

OpenStudy (anonymous):

so it would be x<1670?

OpenStudy (anonymous):

@saranya its not pythagreon therom though

OpenStudy (saranya):

Acc to yu , @ joy If you try to make a triangle with sticks 620mm and 1050mm, you will soon see the 3rd side is always less thanb 620+1050 = 1670mm So the answer is 1670 acc to yu ! Am i ryt ??

OpenStudy (anonymous):

idk if its right though im gonna try it though thanks anyways

OpenStudy (shubhamsrg):

what are you gonna try in that..theres no alternative buddy..ofcorse thats right..

OpenStudy (anonymous):

its for flvs